OpenStudy (anonymous):

Evaluate 9x2y-2 for x = –3 and y = 2.

OpenStudy (anonymous):

put x=-3 where x is in equation and y=2 in equation where y is

OpenStudy (anonymous):

so far what i got is 9-3^22^2

OpenStudy (anonymous):

9(-3)^2(2-2) =0 because 2-2=0

OpenStudy (anonymous):

if question is (9x^2y)-2 then (9*9*2)-2=168-2=166

OpenStudy (anonymous):

whats your question actually

OpenStudy (anonymous):

oops i did mistakeits 2y=4 (9*9*4)-2=324-2=322